# Limits to extreme event forecasting in chaotic systems

### v1

@@ 2 Methods
The minimum probability of error in the extreme event forecaster can be decomposed into three sources, $P^{c}_{e}(\boldsymbol{\hat{Q}}^{-},\hat{f})=P^{c}_{e,I}+P^{c}_{e,O}+P^{c}_{e,M}$ given by
@@
where $\boldsymbol{Q}^{-}=\boldsymbol{Q}(t-\delta t)\pm\delta\boldsymbol{Q}$ contains all the degrees of freedom governing the system (i.e., absolute observability) with finite precision, $P^{c}_{e,I}$ represents the probability of error solely due to uncertainty in the initial conditions, $P^{c}_{e,O}$ denotes the probability of error caused by missing information from unobserved variables, and $P^{c}_{e,M}$ is the probability of error attributable to a suboptimal model.
In the following, we demonstrate the application of our results in two distinct scenarios: the Rössler system and the Kolmogorov flow. The Rössler system offers a simple case for studying extreme events in a chaotic system where all variables can be observed. We use this case to illustrate the classification of errors from Eq. (6). On the other hand, the Kolmogorov flow, characterized by complex, multi-scale interactions among numerous degrees of freedom, represents the dynamics of extreme events found in realistic systems. This case is used to demonstrate the effect of risk-sensitive analysis.
@@ 3.1 Rössler system
with parameters $a=0.1,b=0.1$, and $c=14$. The extreme events are based on $\theta_{3}$, which exhibits rare excursions of intense magnitude. Figure 1(a) shows the trajectory of the Rössler system in the three-dimensional phase space. The extreme event indicator is defined as
@@
where $\bar{\theta}_{3}$ and $\sigma_{\theta_{3}}$ are the mean and standard deviation of ${\theta_{3}}$ over time. Figure 1(b) contains a fragment of the time history of $\boldsymbol{Q}$ and the extreme event indicator $E$. We investigate the case with balanced risk $c^{+}=c^{-}=1$.
We define the normalized probability of error as $\bar{P}_{e}^{c}=P_{e}^{c}/P(E=1)$, where $P(E=1)$ is the probability of extreme events such that $\bar{P}_{e,\min}^{c}\rightarrow 1$ for $\delta t\rightarrow\infty$. Figure 2 shows the normalized minimum probability of error as a function of $\delta t$. Two scenarios are considered. In the first case (represented by a solid line), we assume the only observable variable is $\hat{Q}^{-}=\theta_{3}(t-\delta t)+\delta\theta_{3}$, i.e., $P_{e,\min}^{c}(\hat{Q}^{-})$, where the uncertainty $\delta\theta_{3}$ is 10% of $\sigma_{\theta_{3}}$. In the second scenario (dashed line), it is presumed that knowledge of the full state $\boldsymbol{Q}^{-}$ is available, i.e., $P_{e,\min}^{c}(\boldsymbol{Q}^{-})$. The region beneath each curve corresponds to models that are unattainable, while the region above represents models that are underperforming. The two curves also allow us to discern between errors from uncertainty in initial conditions ($P^{c}_{e,I}$) and errors from the unobserved variables $\theta_{1}$ and $\theta_{2}$ ($P^{c}_{e,O}$). Over time, both cases converge to $\bar{P}_{e}^{c}\rightarrow 1$ given the chaotic nature of the system. This convergence is slower for $P_{e,\min}^{c}(\boldsymbol{Q}^{-})$, as errors are solely due to uncertainties in the initial condition.
@@
Figure 1: (a) Trajectory of the Rössler system. (b) Extraction of time series of $\theta_{1}$, $\theta_{2}$, $\theta_{3}$ and $D$ in the Rössler system. Although not shown, the whole time-span of the signals is 10,000 time units.
@@
Figure 2: Normalized probability of error for extreme events prediction in the Rössler system. $P^{c}_{e,I}$: error due to uncertainty in the initial conditions; $P^{c}_{e,O}$: error caused by unobserved variables; $P^{c}_{e,M}$: error attributable to a suboptimal model.
We also consider a logistic regression model that uses $\theta_{3}$ as input and is trained for different forecast times $\delta t$ (also included in Fig. 2). The probability of error for the logistic regression model, $P_{e}^{c}(\hat{Q}^{-},\hat{f}^{LR})$, enables the quantification of the model error ($P^{c}_{e,M}$) for this specific case. There is a significant gap between $P^{c}_{e,M}$ and the error bounds, implying that potential model improvements are possible. As $\delta t$ increases, $P^{c}_{e,M}$ rapidly becomes the predominant source of error: $P^{c}_{e,I}$ is minor compared to $P^{c}_{e,O}$ and $P^{c}_{e,M}$, indicating that the missing variables and the suboptimal model have a more significant impact on the accuracy of the forecast compared to errors in the initial condition.
@@ 3.2 Kolmogorov flow
$\frac{\partial u_{i}}{\partial t}$ $=-\frac{\partial(u_{i}u_{j})}{\partial x_{j}}-\frac{\partial p}{\partial x_{i}}+\frac{1}{\text{Re}}\frac{\partial^{2}u_{i}}{\partial x_{k}\partial x_{k}}+f_{i},\quad i=1,2,$ (9)
@@
where repeated indices imply summation, $u_{i}(x_{1},x_{2},t)$ is the velocity $i$-th velocity component, $p$ is the pressure, $x_{1}$ and $x_{2}$ are the spatial coordinates, and $f_{i}$ is the forcing with $f_{1}=\sin(4x_{2})$ and $f_{2}=0$. The flow setup is characterized by the Reynolds number $\text{Re}\approx 100$ for which the flow exhibits intermittent bursts of dissipation events Zeff2003 . For our analysis, we use data from Farazmand and Sapsis Farazmand2017 , obtained by numerically resolving all the scales of the problem in a doubly periodic box with side $2\pi$ and $256^{2}$ spatial Fourier modes. Figure 3(a) shows the velocity amplitude $|\boldsymbol{u}(x_{1},x_{2},t)|$ at a given time.
@@
where $\bar{D}$ and $\sigma_{D}$ are the mean and standard deviation of ${D}$ over time. The observable chosen is the magnitude of the spatial Fourier mode of velocity vector $\boldsymbol{u}=[u_{1},u_{2}]$ corresponding to the wavenumber $[1,0]$, which is denoted by $\hat{Q}=\left|\boldsymbol{\breve{u}}_{1,0}\right|$. An excerpt of $\hat{Q}$ and $D$ extracted from the full time history is presented in Fig. 3(b). The time is non-dimensionalized by $t_{e}=(\text{Re}\bar{D})^{-1/2}$.
@@
Figure 3: (a) The instantaneous velocity amplitude $|\boldsymbol{u}|$ for the Kolmogorov flow. (b) Time history of $D(t)$ and $|\boldsymbol{\breve{u}}_{1,0}|(t)$.
@@
Three cases are investigated: $c^{-}$ = 1, 1.5, and 2. The first case ($c^{-}$ = 1) penalizes false positives and false negatives equally. The other two cases assign a higher penalty to false negatives, such that the cost of failing to predict an extreme event is two to three times the cost of incorrectly predicting a non-extreme event. We denote the minimum probability of error for each case as $P_{e,\min}^{c^{-}=1}(\hat{Q}^{-})$, $P_{e,\min}^{c^{-}=1.5}(\hat{Q}^{-})$, and $P_{e,\min}^{c^{-}=2}(\hat{Q}^{-})$, respectively. Figure 4 shows the results as a function of $\delta t$. The minimum cost-sensitive probability of error is normalized as $\bar{P}_{e,\min}^{c}=P_{e,\min}^{c}/\min\{\left(c^{-}\cdot P(E=1),c^{+}\cdot(1-P(E=1)\right)\}$ such that $\bar{P}_{e,\min}^{c}\rightarrow 1$ for $\delta t\rightarrow\infty$.
@@
Figure 4: Normalized, cost-sensitive minimum probability of error for extreme event prediction in the Kolmogorov flow.
@@
The results from Fig. 4 show that increasing $c^{-}$ reduces the normalized minimum probability of error, making the predictions less challenging. This trend is particular to the Komogorov flow and the chosen variables, and other systems can exhibit different behavior. Figure 4 also illustrate two interesting characteristics of $\bar{P}_{e,\min}^{c}(\hat{Q}^{-})$ for the three $c^{-}$ values considered. First, $\bar{P}_{e,\min}^{c}(\hat{Q}^{-})$ is not initially zero when $\delta t=0$. This situation arises because $D$ was not considered an observed variable, leading to uncertainty in $E$ even at $\delta t=0$. The second interesting observation is that $\bar{P}_{e,\min}^{c}$ reaches its lowest value at $\delta t_{\min}=2.8t_{e}$. This observation can be understood by noting that in this system, energy is transferred among different scales due to nonlinear interactions until it is ultimately dissipated. This process occurs on a timescale comparable to $\delta t_{\min}$ (Farazmand2017, ), which explains the effectiveness of $\left|\boldsymbol{\breve{u}}_{1,0}\right|$ in predicting extreme dissipation events at that time lag. For times beyond this point, $\left|\boldsymbol{\breve{u}}_{1,0}\right|$ becomes less useful due to the chaoticity of turbulence.

### v2

@@ 2.4 Sources contributing to the probability of error
The probability of error in the extreme event forecaster can be decomposed into three sources: $P^{c}_{e}(\boldsymbol{\hat{Q}}^{-},\hat{f})=P^{c}_{e,I}+P^{c}_{e,O}+P^{c}_{e,M}$ given by
@@
where the vector $\boldsymbol{Q}^{-}=\boldsymbol{Q}\pm\delta\boldsymbol{Q}$ contains all the degrees of freedom governing the system $\boldsymbol{Q}$ (i.e., absolute observability) but with finite precision $\delta\boldsymbol{Q}$. The interpretation of each term in Eq. (13) is as follows:
    $P^{c}_{e,I}$ represents the probability of error solely arising from uncertainty in the initial conditions. This is because $\boldsymbol{Q}^{-}$ contains all the degrees of freedom of the system, which are sufficient to integrate the system forward in time. However, the process is subject to the initial uncertainty $\delta\boldsymbol{Q}$ such that a higher $\delta\boldsymbol{Q}$ might result in a higher $P^{c}_{e,I}$. The magnitude of $\delta\boldsymbol{Q}$ varies depending on the specific problem. If $\boldsymbol{Q}^{-}$ is known with infinite precision (i.e., $\delta\boldsymbol{Q}=\boldsymbol{0}$), then $P^{c}_{e,I}$ equals zero.
    $P^{c}_{e,O}$ denotes the probability of error caused by missing information from unobserved variables. This error originates from the fact that $\boldsymbol{\hat{Q}}^{-}$ contains less information than $\boldsymbol{Q}^{-}$. As discussed in §2.3, the inclusion of multiple time lags in $\boldsymbol{\hat{Q}}^{-}$ can compensate for the lack of observed variables Takens1981 . However, $P^{c}_{e,O}$ will still be an important contributor to the total probability of error in those situations where the number of degrees of freedom is much larger than the number of observed variables, $N\gg M$.
    $P^{c}_{e,M}$ is the probability of error attributable to a suboptimal model. Values of $P^{c}_{e,M}>0$ imply that $\hat{f}$ is not efficiently exploiting the information available in $\boldsymbol{\hat{Q}}^{-}$. In those cases, the model is not operating at its theoretical maximum performance, and further improvements are possible. Conversely, $P^{c}_{e,M}=0$ implies that $\hat{f}$ is the best-performing model given the observed variables and uncertainties in the initial conditions.
In the following, we demonstrate the application of our results in two distinct scenarios: the Rössler system and the Kolmogorov flow. The Rössler system offers a simple case for studying extreme events in a chaotic system where all variables can be observed. We use this case to illustrate the classification of errors from Eq. (13). On the other hand, the Kolmogorov flow, characterized by complex, multi-scale interactions among numerous degrees of freedom, represents the dynamics of extreme events found in more realistic systems. This case is used to demonstrate the effect of cost-sensitive analysis.
@@ 3.1 Rössler system
with parameters $a=0.1,b=0.1$, and $c=14$. We investigate extreme events in $\theta_{3}$, which exhibits rare excursions of intense magnitude. Figure 1(a) shows the trajectory of the Rössler system in the three-dimensional phase space. The extreme event indicator is defined as
@@
where $\bar{\theta}_{3}$ and $\sigma_{\theta_{3}}$ are the mean and standard deviation of ${\theta_{3}}$ over time. The threshold is set to $\eta=\bar{\theta}_{3}+3\sigma_{\theta_{3}}$, but the conclusions drawn in this section apply to other values of $\eta$. Results for a higher threshold can be found in D. Figure 1(b) contains a fragment of the time history of $\boldsymbol{Q}$ and the extreme event indicator $E$.
We investigate the case with balanced risk $c^{+}=c^{-}=1$ and define the normalized probability of error as $\bar{P}_{e}^{c}=P_{e}^{c}/C$. This normalization is such that $\bar{P}_{e,\min}^{c}\rightarrow 1$ for $\delta t\rightarrow\infty$ in practical applications. Figure 2 shows the normalized minimum probability of error as a function of time-horizon for extreme event prediction $\delta t$ using Eq. (7). Three scenarios are considered.
    In the first case, we assume that the only observable variable is $\hat{Q}^{-}_{1}=\theta_{3}(t)\pm\delta\theta_{3}(t)$, where the uncertainty in the initial condition is set to $\delta\theta_{3}=0.05\sigma_{\theta_{3}}$. Here, the uncertainty $\delta\theta_{3}$ is not introduced by perturbing the equations of the system. Instead, the uncertainty is incorporated in a non-intrusive manner when calculating the probability $P(\theta_{3})$ by discretizing $\theta_{3}$ into bins of size $2\delta\theta_{3}$. This is equivalent to assuming that the solution passing through $\theta_{3}$ cannot be distinguished from another trajectory, also contained within the attractor of the system, at a distance from $\theta_{3}$ equal to or less than $\delta\theta_{3}$. The associated minimum probability of error, $\bar{P}_{e,\min}^{c}(\hat{Q}^{-}_{1})$, is represented by the solid line in Fig. 2(a).
    In the second scenario, the observable includes two time lags in addition to the present time: $\boldsymbol{\hat{Q}}^{-}_{2}=[\theta_{3}(t),\theta_{3}(t-\delta t),\theta_{3}(t-2\delta t)]\pm\delta\boldsymbol{\theta}_{3}$, where the uncertainty $\delta\boldsymbol{\theta}_{3}$ is again set to $0.05\sigma_{\theta_{3}}$ for all time lags. The minimum probability of error, $\bar{P}_{e,\min}^{c}(\boldsymbol{\hat{Q}}^{-}_{2})$, is depicted by the solid line in Fig. 2(b). The difference between the two minimal errors $\bar{P}_{e,\min}^{c}(\boldsymbol{\hat{Q}}^{-}_{2})$ and $\bar{P}_{e,\min}^{c}(\hat{Q}^{-}_{1})$ from Fig. 2(a) serves as a measure of the improvement in predictive accuracy gained by incorporating observations from two additional times in $\theta_{3}$.
    In the third scenario, it is assumed that the knowledge of the full state is available at the present time with finite precision, i.e., $\boldsymbol{Q}^{-}=[\theta_{1}(t),\theta_{2}(t),\theta_{3}(t)]\pm\delta\boldsymbol{\theta}$, where $\delta\boldsymbol{\theta}=0.05[\sigma_{1},\sigma_{2},\sigma_{3}]$ with $\sigma_{i}$ the standard deviation of $\theta_{i}$. The minimum probability of error, $\bar{P}_{e,\min}^{c}(\boldsymbol{Q}^{-})$, is indicated by the dashed line in Fig. 2(a) and Fig. 2(b). Errors arising from uncertainty in initial conditions are quantified by $\bar{P}^{c}_{e,I}=\bar{P}_{e,\min}^{c}(\boldsymbol{Q}^{-})$ (highlighted by the purple shaded region in Fig. 2). The discrepancy between $\bar{P}_{e,\min}^{c}(\hat{Q}^{-}_{1})$ and $\bar{P}_{e,\min}^{c}(\boldsymbol{Q}^{-})$ in Fig. 2(a), and between $\bar{P}_{e,\min}^{c}(\boldsymbol{\hat{Q}}^{-}_{2})$ and $\bar{P}_{e,\min}^{c}(\boldsymbol{Q}^{-})$ in Fig. 2(b), allows us to quantify the errors resulting from the lack of knowledge of $\theta_{1}$ and $\theta_{2}$ (i.e., $\bar{P}^{c}_{e,O}$, indicated by the yellow shaded region in Fig. 2).
The region beneath each curve, $\bar{P}_{e,\min}^{c}(\hat{Q}^{-}_{1})$, $\bar{P}_{e,\min}^{c}(\boldsymbol{\hat{Q}}^{-}_{2})$, and $\bar{P}_{e,\min}^{c}(\boldsymbol{Q}^{-})$, corresponds to models that are unattainable given the observable $\hat{Q}^{-}_{1}$, $\boldsymbol{\hat{Q}}^{-}_{2}$, and $\boldsymbol{Q}^{-}$, respectively, whereas the region above represents models that are suboptimal. Over time, all cases converge to $\bar{P}_{e}^{c}\rightarrow 1$ given the chaotic nature of the system. This convergence is slower for $\bar{P}_{e,\min}^{c}(\boldsymbol{Q}^{-})$, as errors are only due to uncertainties in the initial condition.
To illustrate the errors from an actual predictive model, we trained decision tree models, $\hat{f}^{DT}$, to predict $E$ using either $\hat{Q}^{-}_{1}$ or $\boldsymbol{\hat{Q}}^{-}_{2}$ as input. Different decision tree models are trained to forecast $E$ at each $\delta t$. The maximum number of branch node splits is 8, and each leaf contains at least 10 observations. The results are also included in Fig. 2(a) and (b). Additional details about the confusion matrix for the decision tree model can be found in D. The normalized probability of error for the decision tree model, Both $\bar{P}_{e}^{c}(\hat{Q}^{-}_{1},\hat{f}^{DT})$ and $\bar{P}_{e}^{c}(\boldsymbol{\hat{Q}}^{-}_{2},\hat{f}^{DT})$ enable the quantification of the model error $\bar{P}^{c}_{e,M}$ for the specific case. The results in Fig. 2(a) show that the model error $\bar{P}_{e}^{c}(\hat{Q}^{-}_{1},\hat{f}^{DT})$ closely approaches the minimum error given by $\bar{P}_{e,\min}^{c}(\hat{Q}^{-}_{1})$, indicating that the model is operating near its maximum theoretical performance. On the other hand, the results in Fig. 2(b) reveal a gap between the model error $\bar{P}_{e}^{c}(\boldsymbol{\hat{Q}}^{-}_{2},\hat{f}^{DT})$ and the minimum theoretical error $\bar{P}_{e,\min}^{c}(\boldsymbol{\hat{Q}}^{-}_{2})$ for $\delta t<0.4$. This indicates that the model is suboptimal, and models with improved performance are possible. For both cases, as $\delta t$ increases, $\bar{P}^{c}_{e,O}$ rapidly becomes the predominant source of error. Conversely, $\bar{P}^{c}_{e,I}$ is minor compared to $\bar{P}^{c}_{e,O}$. Hence, the analysis also shows that missing variables have a greater impact on the accuracy of the forecast compared to suboptimal modeling and uncertainty in the initial conditions.
@@
Figure 1: (a) Trajectory of the Rössler system. (b) Extraction of time series of $\theta_{1}$, $\theta_{2}$, $\theta_{3}$ and extreme events in the Rössler system. Although not shown, the whole time-span of the signals is 10,000 time units.
@@
Figure 2: Normalized probability of error for extreme events prediction in the Rössler system for the threshold $\eta=\bar{\theta}_{3}+3\sigma_{\theta_{3}}$ using as observable (a) $\hat{Q}^{-}_{1}=\theta_{3}(t)\pm\delta\theta_{3}$ and (b) $\boldsymbol{\hat{Q}}^{-}_{2}=[\theta_{3}(t),\theta_{3}(t-\delta t),\theta_{3}(t-2\delta t)]\pm\delta\boldsymbol{\theta}_{3}$ $\bar{P}_{e,\min}^{c}(\hat{Q}^{-}_{1})$, $\bar{P}_{e,\min}^{c}(\boldsymbol{\hat{Q}}^{-}_{2})$, and $\bar{P}_{e,\min}^{c}(\boldsymbol{Q}^{-})$ are the minimum probability of error using the observable $\hat{Q}^{-}_{1}$, $\boldsymbol{\hat{Q}}^{-}_{2}$, and $\boldsymbol{Q}^{-}=[\theta_{1}(t),\theta_{2}(t),\theta_{3}(t)]\pm\delta\boldsymbol{\theta}$, respectively. $\bar{P}^{c}_{e,I}$ (purple) is the error due to uncertainty in the initial conditions; $\bar{P}^{c}_{e,O}$ (yellow) is the error caused by unobserved variables; $\bar{P}^{c}_{e,M}$ (red) is the error due to suboptimal model.
Finally, we compare the exact minimum probability of error from Eq. (7) with the information-theoretic bounds from Eq. (11) for the second-order, cost-sensitive conditional Rényi entropy. The results are presented in Fig. (3) using either $\hat{Q}^{-}_{1}$ or $\boldsymbol{\hat{Q}}^{-}_{2}$ as observables. In both cases, the bounds provide a narrow region within which $\bar{P}_{e,\min}^{c}$ must be confined. In situations where directly obtaining $\bar{P}_{e,\min}^{c}$ is challenging, the region defined by the upper and lower bounds can be used to demarcate the theoretical zone of near-optimal operation for a model. If the error falls within this zone, the model can be considered as possibly operating near its best theoretical performance.
Figure 3: Normalized information-theoretic upper and lower bounds of the minimum probability of error for extreme events prediction in the Rössler system for the threshold $\eta=\bar{\theta}_{3}+3\sigma_{\theta_{3}}$ and observables $\hat{Q}^{-}_{1}$ and $\boldsymbol{\hat{Q}}^{-}_{2}$. The solid line is $\bar{P}_{e,\min}^{c}$ and the shaded area represents the region confined within $\bar{P}_{e,\text{min,LB}}$ and $\bar{P}_{e,\text{min,UB}}$ obtained for the second-order, cost-sensitive conditional Rényi entropy.
@@ 3.2 Kolmogorov flow
$\frac{\partial u_{i}}{\partial t}$ $=-\frac{\partial(u_{i}u_{j})}{\partial x_{j}}-\frac{\partial\Pi}{\partial x_{i}}+\frac{1}{\text{Re}}\frac{\partial^{2}u_{i}}{\partial x_{k}\partial x_{k}}+f_{i},\quad i=1,2,$ (16)
@@
where repeated indices imply summation, $u_{i}(x_{1},x_{2},t)$ is the $i$-th velocity component, $\Pi$ is the pressure, $x_{1}$ and $x_{2}$ are the spatial coordinates, and $f_{i}$ is the forcing with $f_{1}=\sin(4x_{2})$ and $f_{2}=0$. The velocity vector is denoted as $\boldsymbol{u}=[u_{1},u_{2}]$. The flow setup is characterized by the Reynolds number $\text{Re}=100$ for which the flow exhibits intermittent bursts of dissipation events Zeff2003 . For our analysis, we use data from Farazmand and Sapsis Farazmand2017 , obtained by numerically resolving all the scales of the problem in a doubly periodic box with side $2\pi$ and $256^{2}$ spatial Fourier modes. Figure 4(a) shows the velocity amplitude $|\boldsymbol{u}(x_{1},x_{2},t)|$ at a given time.
@@
where $\bar{D}$ and $\sigma_{D}$ are the mean and standard deviation of ${D}$ over time. Predictions of extreme events using a higher threshold can be found in B. The observable chosen is the magnitude of the spatial Fourier mode of $\boldsymbol{u}$ corresponding to the wavenumber $[1,0]$, which is denoted by $\hat{Q}=\left|\boldsymbol{\breve{u}}_{1,0}\right|$. The latter is one of the preferred observables for predicting extreme dissipation events in the Kolmogorov flow, as it has been shown to correlate with the growth of $D$ Farazmand2017 . An excerpt of $\hat{Q}$ and $D$ extracted from the full time history is presented in Fig. 4(b). The time is non-dimensionalized by $t_{e}=(\text{Re}\bar{D})^{-1/2}$.
@@
Figure 4: (a) The instantaneous velocity amplitude $|\boldsymbol{u}|$ for the Kolmogorov flow. (b) Time history of $D(t)$ and $|\boldsymbol{\breve{u}}_{1,0}|(t)$ for $\eta$ = 1.5.
@@
Three cases are investigated with increasing cost for false positives: $c^{-}$ = 1, 1.5, and 2. The first case ($c^{-}$ = 1) penalizes false positives and false negatives equally. The other two cases assign a higher penalty to false negatives, such that the cost of failing to predict an extreme event is two times (for $c^{-}=1.5$) or three times (for $c^{-}=2$) the cost of incorrectly predicting a non-extreme event. We denote the minimum probability of error for each case as $P_{e,\min}^{c^{-}=1}(\hat{Q}^{-})$, $P_{e,\min}^{c^{-}=1.5}(\hat{Q}^{-})$, and $P_{e,\min}^{c^{-}=2}(\hat{Q}^{-})$, respectively. Figure 5 shows the results as a function of $\delta t$. The minimum cost-sensitive probability of error is normalized as $\bar{P}_{e,\min}^{c}=P_{e,\min}^{c}/C$ such that $\bar{P}_{e,\min}^{c}\rightarrow 1$ for $\delta t\rightarrow\infty$.
@@
Figure 5: Normalized, cost-sensitive, minimum probability of error for extreme events prediction in the Kolmogorov flow for $\eta=\bar{D}+1.5\sigma_{D}$ and $c^{-}=1$, 1.5, and 2.
@@
The results from Fig. 5 show that increasing $c^{-}$ reduces the normalized minimum probability of error, making the predictions less challenging. This trend is particular to the Komogorov flow and the chosen variables, and other systems can exhibit different behavior. Figure 5 also illustrates two interesting characteristics of $\bar{P}_{e,\min}^{c}(\hat{Q}^{-})$ for the three $c^{-}$ values considered. First, $\bar{P}_{e,\min}^{c}(\hat{Q}^{-})$ is not initially zero when $\delta t=0$. This situation arises because $D$ was not considered an observed variable, leading to uncertainty in $E$ even at $\delta t=0$. The second interesting observation is that $\bar{P}_{e,\min}^{c}$ reaches its lowest value at $\delta t_{\min}=2.8t_{e}$. This observation can be understood by noting that in this system, energy is transferred among different scales due to nonlinear interactions until it is ultimately dissipated. This process occurs on a timescale comparable to $\delta t_{\min}$ (Farazmand2017, ), which explains the effectiveness of $\left|\boldsymbol{\breve{u}}_{1,0}\right|$ in predicting extreme dissipation events at that time lag. For times beyond this point, $\left|\boldsymbol{\breve{u}}_{1,0}\right|$ becomes increasingly less effective due to the chaoticity of turbulence.
Finally, we compare the normalized information-theoretic upper and lower bounds of the minimum probability of error when $c^{-}=2$. The results, presented in Fig. 6, show that the bounds accurately reflect the trend observed for $\bar{P}_{e,\min}^{c^{-}=2}$: there is a non-zero minimum probability of error at $\delta t=0$, which initially decreases before eventually increasing towards one.
Figure 6: Normalized, cost-sensitive, information-theoretic upper and lower bounds of the minimum probability of error for extreme events prediction in the Kolmogorov flow for $\eta=\bar{D}+1.5\sigma_{D}$. The solid line is $\bar{P}_{e,\min}^{c}$ and the shaded area represents the region confined within $\bar{P}_{e,\text{min,LB}}$ and $\bar{P}_{e,\text{min,UB}}$ obtained for the cost-sensitive, second-order, conditional Rényi entropy.